A bullet with a mass of 4 grams (g) is fired from a gun of mass 5kg with a velocity of 600m/s-1 What is the velocity of recoil of the gun?


1
Expert's answer
2021-04-20T16:49:17-0400

"\text{bullet mass and velocity}"

"m_b= 8*10^{-3}kg;V_{b1}= 600m\/s;V_{b0}=0;"

"\text{gun mass and speed:}"

"m_g = 5 kg;V_{g1};V_{g0}=0;"

"\text{according to the law of conservation of momentum:}"

"m_b\Delta V_b+m_g\Delta V_g=0"

"\Delta V_g=-\frac{m_b\Delta V_b}{m_g}=-\frac{8*10^{-3}*600}{5}=-0.96m/s"

"V_{g1}= -0.96m\/s"

"\text{}" "\text{minus sign means that the direction of movement}\newline
\text{ of the bullet and the gun are opposite}"

Answer:0.96 m/s velocity of recoil of the gun